Question:A 2 kg sphere falls from a height 3 m.If whole of its potential energy is converted into heat ,then how much heat will be produced ?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Heat = m*g*h
 = 2 *9.8* 3 = 58.8 Joules = 58.8/4.2 Calories
 = 14 Calori
 
Regards
Arun (askIITians forum expert)